### v2

# Distillation Robustifies Unlearning

Abstract
Current LLM unlearning methods are not robust: they can be reverted easily with a few steps of finetuning. This is true even for the idealized unlearning method of training to imitate an oracle model that was never exposed to unwanted information, suggesting that output-based finetuning is insufficient to achieve robust unlearning. In a similar vein, we find that training a randomly initialized student to imitate an unlearned model transfers desired behaviors while leaving undesired capabilities behind. In other words, distillation robustifies unlearning. Building on this insight, we propose Unlearn-Noise-Distill-on-Outputs (UNDO), a scalable method that distills an unlearned model into a partially noised copy of itself. UNDO introduces a tunable tradeoff between compute cost and robustness, establishing a new Pareto frontier on synthetic language and arithmetic tasks. At its strongest setting, UNDO matches the robustness of a model retrained from scratch with perfect data filtering while using only 60-80% of the compute and requiring only 0.01% of the pretraining data to be labeled. We also show that UNDO robustifies unlearning on the more realistic Weapons of Mass Destruction Proxy (WMDP) benchmark. Since distillation is widely used in practice, incorporating an unlearning step beforehand offers a convenient path to robust capability removal.
33footnotetext: We share our code implementation publicly through GitHub.
1 Introduction
Large language models (LLMs) can acquire harmful capabilities during pretraining on massive, largely unfiltered datasets [9, 62]. This complicates model deployment. For example, a model with knowledge relevant to developing novel cyberweapons could enable global-scale harm if accessed by bad actors [52, 61]. While data filtering before pretraining could mitigate such risks, precisely auditing and filtering data at the required scale remains impractical [54, 81, 1].
Post-training methods like Reinforcement learning from Human Feedback discourage models from using undesired capabilities on common inputs [4, 53], but leave the underlying capabilities intact. As a result, the model remains vulnerable to attacks that can elicit these capabilities, including adversarial prompts, jailbreaks, or direct finetuning [83, 36, 80]. To address this vulnerability, recent work has tried using machine unlearning to remove undesired capabilities from LLMs [51, 66, 6]. However, existing unlearning methods also suppress capabilities rather than remove them [30, 18]. The supposedly unlearned capabilities can be restored through a few steps of finetuning, leaving the fundamental challenge of achieving true capability removal unsolved [46, 47, 22, 31].
We present a simple but powerful observation: distillation robustifies unlearning. When an unlearned model is distilled into a randomly initialized student, desired behavior is transferred while undesired capabilities are left behind (Figure 1). Across diverse domains, including language, arithmetic, and weapons of mass destruction, we find that distillation consistently improves resistance to adversarial relearning attacks. Remarkably, distilled models relearn harmful capabilities at rates comparable to an oracle model [24] trained from scratch with data filtering.
This finding has immediate practical implications. Distillation is already standard practice in many contexts [29, 82]. For example, LLM developers distill models to reduce inference costs [71, 72]. By applying unlearning methods before distillation, these developers can produce models that are strongly resistant to relearning attacks.
Building on this insight, we introduce UNDO (Unlearn-Noise-Distill-on-Outputs), which distills an unlearned model into a partially noised copy of itself, enabling a tradeoff between compute cost and unlearning robustness. With this framework, we approximate the robustness of gold-standard data filtering at a fraction of the cost, offering a new frontier in safe and scalable model deployment. We summarize our contributions as follows:
  • In Section 3, we provide evidence of a limitation in behavioral training for robust unlearning in LLMs. Even when achieving perfect behavioral agreement with an unlearning oracle [24], models retain latent capabilities that can be easily restored through finetuning.
  • In Section 4, we show that distillation robustifies unlearning. Training a new, randomly initialized model on the outputs of an unlearned model will transfer desired behavior while leaving latent capabilities behind.
  • In Section 5, we propose UNDO, which distills an unlearned model’s outputs into a noised copy of itself. This method enables a novel compute-robustness tradeoff that interpolates between mere behavioral suppression and fully robust unlearning.
  • We benchmark UNDO against a variety of unlearning methods on language and arithmetic tasks, extending the Pareto frontier of retain performance versus forget unlearning robustness. We also show competitive performance and robustification of unlearning on the WMDP benchmark [40].
Figure 1: Distillation robustifies unlearning. Existing LLM unlearning methods suppress undesired behavior, but are reversible using a small amount of finetuning. We show that distilling the suppressed model into a randomly initialized network significantly increases resilience against reacquiring the undesired behavior. Our method substantially outperforms other robust unlearning baselines, including RepNoise [59] and SAM [22].
2 Robust Unlearning
Machine unlearning seeks to remove specified knowledge from a trained model while preserving the model’s overall functionality. This goal has been operationalized in different ways [39]. We follow recent work concerned with removing undesired capabilities from LLMs in a way that is robust to adversarial elicitation [65, 68]. We consider the following problem:
Robust Unlearning. A robust unlearning method accepts the following inputs: (i) a Reference Model: a neural network with capabilities to be removed; (ii) a Retain Set: a collection of examples demonstrating desired capabilities; (iii) a Forget Set: a collection of examples demonstrating undesired capabilities; (iv) Training Data: a collection of unlabeled examples. The method produces a new model that is evaluated according to (a) a Retain Evaluation: a measure of the model’s desired capabilities; (b) an Elicited Forget Evaluation: a measure of the model’s latent undesired capabilities. The goal is to produce a model with a high retain performance and low elicited forget performance.
Elicitation Mechanism. We use finetuning on the forget set, commonly referred to as a relearning attack, as our elicitation method. Relearning attacks have repeatedly proven to be among the most effective ways to elicit unlearned knowledge or capabilities in supposedly unlearned models [31, 58]. Therefore, recent studies treat resistance to such finetuning-based probes as the primary criterion for robust unlearning [11, 22, 16]. For rigorous evaluation, we apply several relearning attacks at different learning rates and measure the maximum elicited forget performance of all attacks.
Reference vs. Oracle Models. In our framework, the reference model represents a trained model that contains both desired and undesired capabilities (e.g., a pretrained LLM). We also define an oracle model as one trained on the reference model’s pretraining dataset with all content directly related to the undesirable capability perfectly filtered out. The oracle model often represents a gold standard solution to the robust unlearning problem, simulating perfect capability removal by never learning undesired capabilities in the first place [48, 84].
The capabilities-focused definition used above differs from machine unlearning used in privacy and fairness applications [28, 86]. Machine unlearning in privacy and fairness applications assumes the retain and forget sets partition the reference model’s training data, and aims to produce an oracle model equivalent to training solely on the retain set [24]. In contrast, the capabilities-focused definition uses retain and forget sets as curated datasets reflecting desired and undesired capabilities, which need not correspond directly to literal subsets of the training data [13, 17]. Despite this conceptual shift, similar methods are often applied across both settings; however, these methods generally do not achieve robustness [75, 32].
3 Oracle Matching Does Not Guarantee Robust Unlearning
Unlearning methods often finetune the outputs of the reference model. For example, Gradient Ascent combined with Gradient Descent maximizes cross-entropy loss on the forget set and minimizes it on the retain set [48]. The aspirational goal of such procedures is oracle matching: attaining outputs indistinguishable from those of a model that never encountered the undesired data.
In this section, we demonstrate the limitations of this aspiration empirically. We finetune a reference model to match an oracle model’s outputs on both the retain and forget sets to achieve effectively perfect behavioral agreement. Yet, when both models are subsequently finetuned on the forget set, the oracle-matched reference model relearns significantly faster than the oracle model. This shows that optimizing for behavioral outputs is likely insufficient for robust unlearning.
Datasets and Models. We conduct experiments in a language setting and an arithmetic setting, defined below. For each setting, we train a reference model with both retain and forget capabilities, and an oracle model with only retain capabilities by construction.
More details on how these datasets and reference models were generated can be found in Appendix A.
  • Language. The retain set is English documents from FineWebEdu [56]; the forget set is Korean documents from FineWeb2 [57]; the training data is the union of these two sets. The reference model is a 100M parameter model based on the Gemma 2 architecture (see Table 1 for details) pretrained on 2B tokens, 1B from retain data and 1B from forget; the oracle model is the same 100M model architecture pretrained on 1B tokens of the retain set. The retain and forget evaluations are cross-entropy loss on held-out portions of the retain and forget sets, respectively, each containing 500K tokens.
  • Arithmetic. The retain set is addition/subtraction statements (equations and word problems) from a synthetic arithmetic dataset described in Appendix A; the forget set is multiplication and division from the same dataset; the training dataset is the union of these with (English) documents from FineWebEdu [56]; each containing 500K arithmetic questions for training. The reference model is trained for 1000 steps on the training data; the oracle model is trained for 1000 steps on the training data with the forget set replaced with additional English Fineweb data. The retain and forget evaluations measure accuracy on addition/subtraction and multiplication/division problems, using held-out portions of the retain and forget sets.
Oracle Matching Training. We match the reference model’s outputs with those of the oracle model using the Kullback-Leibler Divergence (KL Divergence) as the loss function. For comparison, we also match a randomly initialized model to the oracle model in the same way. The hyperparameters of the distillation setup can be found in Table 3 and the loss curves are shown in Figure 2 (a), where we use the label Student (Reference) or Student (Random) depending on whether the student in the distillation setup is the reference model or the randomly initialized model, respectively. In both settings, the distillation loss tends to zero.
Relearning. After distillation, we apply relearning attacks to all three models, Student (Reference), Student (Random), and Oracle Teacher, for 500 steps on data from the forget set. The hyperparameters for this training are in Table 4. We show the corresponding learning and accuracy curves for multiple learning rates in Figure 2 (b) and (c). The bolded lines correspond to the worst-case adversary (maximum elicitation) for that number of training steps. In both the arithmetic and language settings, Student (Reference) quickly learns the forget distribution compared to the Student (Random) and Oracle Teacher models, which are more robust. This pattern holds both for the worst-case adversary and uniformly across all learning rates.
Figure 2: Matching oracle behavior doesn’t guarantee robust unlearning. (a) KL divergence during distillation shows behavioral alignment with Oracle Teacher. (b-c) Despite this alignment, reference models matched to the oracle (Student (Reference)) exhibit rapid relearning of undesired capabilities when finetuned on the forget set, compared to the randomly initialized model matched to the oracle (Student (Random)) and the oracle itself (Oracle Teacher). Results highlight that an ideal unlearned behavior on the surface is insufficient for ensuring robustness against relearning.
As illustrated in Figure 2, distilling the oracle’s outputs into a randomly initialized student (Student (Random)) results in a model whose relearning speed on the forget distribution closely matches that of the oracle itself. Given that existing machine unlearning methods provide practical means of approximating oracle outputs, an important next question is whether this robustness to relearning remains after replacing true oracle outputs with approximations. In the next section, we demonstrate empirically that they do.
4 Distillation Robustifies Unlearning
The previous section showed that training the reference model to match oracle behavior is insufficient for robust unlearning. However, training a randomly initialized model to match the same oracle outputs does produce a robust unlearning effect. Here, we investigate whether this robustness persists when replacing the oracle with approximations from standard unlearning methods.
Unlearning Methods. We experiment with three unlearning methods from the literature: Gradient Difference (GradDiff) [48], Maximizing Entropy (MaxEnt) [42], and Representation Misdirection for Unlearning (RMU) [40]. These methods represent three major paradigms in LLM unlearning: gradient-based manipulation, output distribution matching, and representation-level interventions.
Unlearning Method A: Gradient Difference (GradDiff) applies opposing gradient updates to forget and retain data. Given a model parameterized by $\theta$, GradDiff minimizes the objective $\mathcal{L}(\theta)\coloneq-L_{\text{forget}}(\theta)+L_{\text{retain}}(\theta)$, where both loss terms are cross-entropy losses. The gradient updates aim to maximize loss on forget examples (pushing probability mass away from undesired outputs) while minimizing loss on retain examples (preserving desired model behavior).
Unlearning Method B: Maximizing Entropy (MaxEnt) increases uncertainty in the model’s outputs on forget data while preserving performance on retain data. MaxEnt minimizes $\mathcal{L}(\theta)\coloneq L_{\text{uniform}}(\theta)+L_{\text{retain}}(\theta)$, where $L_{\text{uniform}}$ is a KL divergence term that pushes the model’s output distribution toward uniformity on forget examples, and $L_{\text{retain}}$ is the standard cross-entropy loss on retain examples. Rather than directly opposing learned associations as in GradDiff, MaxEnt makes the model outputs maximally uninformative on undesired tasks.
Unlearning Method C: Representation Misdirection for Unlearning (RMU) operates at the representation level rather than the output level. It minimizes $\mathcal{L}(\theta)\coloneq L_{\text{misdirect}}(\theta)+L_{\text{preserve}}(\theta)$, where $L_{\text{misdirect}}$ uses mean squared error (MSE) loss to push the model’s internal activations at specific layers toward a random direction for forget examples, and $L_{\text{preserve}}$ applies MSE to maintain similarity to the original model’s activations for retain examples.
Figure 3: Comparing unlearning methods. (a–c) Unlearning trends across hyperparameters for our language setup, where we select configurations that maximize retain performance while minimizing forget performance for distillation (see Figures 4 and 5). (d–f) Corresponding trends in arithmetic.
Experimental Setup. We investigate whether distillation can enhance the robustness of unlearning methods against relearning attacks. Our experimental protocol consists of two phases: (i) applying unlearning methods (GradDiff, MaxEnt, or RMU) to a pretrained model (as shown in Figure 3), and (ii) distilling this suppressed model into a randomly initialized model of identical architecture using forward KL divergence. We refer to this method as Unlearn-and-Distill.
We observe relearning speed when models are subjected to relearning attacks, comparing unlearning-only models, Unlearn-and-Distill models, and oracle baselines. We defer the discussion of other relevant experimental details to Appendix B.
Result: Distillation makes unlearning more resilient to relearning. Figure 4 presents the relearning curves for our experiments across both language and arithmetic domains. The plots track how quickly each model reacquires the forget capability when subjected to relearning attacks. We observe that models that underwent unlearning followed by distillation (Unlearn-and-Distill) are significantly more resistant to relearning compared to unlearning-only counterparts (Unlearn). Notably, in some cases, the distilled models’ relearning trajectories closely approximate those of the gold standard (Data Filtering). This robustness improvement holds whether we measure average performance across learning rates (Figure 10) or worst-case adversarial performance (Figure 4).
While RMU combined with distillation shows less impressive results in the arithmetic domain in Figure 4 (f), it still offers marked improvement over the unlearning-only model. We observe that such underperformance can occur when the initial retain-forget trade-off established by the unlearning method is less favorable. In Figure 3, RMU achieved only 62% retain and 6.8% forget accuracy, compared to MaxEnt’s superior 80% and 1.3% (Section B.1). This suggests that distillation effectiveness can depend on the quality of the teacher model’s retain-forget tradeoff.
Figure 4: Unlearn-and-Distill boosts robustness to relearning. (a-c) Relearning trends for the language forget domain (Korean), comparing unlearning-only methods (GradDiff, MaxEnt, RMU) against models with an additional distillation step, measured against the gold standard of full retraining. We highlight the least favorable learning curve for each method. (d-f) Relearning trends for the arithmetic forget domain (Multiplication & Division).
Why This Matters. To contextualize the experiment, we began with custom-pretrained language models containing dual capabilities. For the language setup, approximately 50% of the pretraining data contributed to desired (retain) capabilities and 50% to undesired (forget) capabilities. Standard unlearning methods use minimal labeled data (less than 0.01% of pretraining data) to suppress undesired capabilities almost perfectly. However, this shallow suppression is vulnerable to relearning attacks. An alternative solution would be data filtering with full retraining, but this approach requires labeling all pretraining data, which is often infeasible at scale. Unlearn-and-Distill bridges this gap by transferring the suppressed model’s behavior to a fresh parameter space using unlabeled pretraining data for distillation, achieving substantial robustness improvements without the extensive labeling requirements of data filtering.
These results provide empirical evidence validating claims of prior work about the challenges of robust unlearning. Some have argued that quick relearning of supposedly unlearned capabilities stems from latent information preserved in the parameter space [88, 55, 64] due to properties like plasticity [45] and representational entanglement [89, 48]. Our experiments confirm that when we apply standard unlearning methods to a model, the resulting model relearns quickly under finetuning. However, when we distill this unlearned model into a randomly initialized student, the student relearns substantially slower while maintaining the same unlearning behavior.
5 Trading Compute for Unlearning Robustness
So far, we have demonstrated that distilling an unlearned model into a randomly initialized model significantly enhances robustness against relearning attacks. However, this approach introduces substantial computational costs, as we are training a student model from scratch. Moving beyond the basic Unlearn-and-Distill method established in Section 4, we now investigate whether the method can be adapted to achieve different levels of robustness with different amounts of compute. That is, can we trade off a bit of robustness to reduce compute costs, or vice versa?
In this section, we conduct experiments that generalize the Unlearn-and-Distill approach in Section 4 with a three-step process. (i) Unlearn: create a behaviorally suppressed model (teacher) by applying standard unlearning methods. (ii) Noise: create a student model by perturbing the weights of the suppressed model, damaging it. (iii) Distill: repair this damaged student by distilling to recover the teacher’s original behavior. The controlled perturbation enables us to interpolate between the original parameters and full reinitialization. We refer to this method as Unlearn-Noise-Distill-on-Outputs, or simply, UNDO. We define the perturbed model as:
$\theta_{\text{perturbed}}=(1-\alpha)\theta_{\text{suppressed}}+\alpha\beta N$
where mixing coefficient $\alpha\in[0,1]$, noise scale $\beta\in\mathbb{R}^{+}$, and $N$ represents sampled noise. We sample $N$ using Xavier initialization, though other random noise sampling methods could be explored. Intuitively, the term $(1-\alpha)\theta_{\text{suppressed}}$ shrinks [2] the suppressed model’s parameters, and the term $\alpha\beta N$ injects noise. As we vary $\alpha$ from 0 to 1, we can view $\theta_{\text{perturbed}}$ as interpolating between the original paramterization and full reinitialization. We observe later in Section 6 that simply shrinking the parameters (setting $\beta=0$) can also be effective, as the key idea is to globally damage the network to varying degrees, though including $\beta$ adds expressivity to the formulation. Note that when $\alpha=1$ and $\beta=1$, the perturbation reduces to random initialization.
Experimental Setup. We start by perturbing the models that are suppressed with MaxEnt from Section 4 and follow the same protocols for distillation using forward KL divergence on the outputs. For both language and arithmetic settings, we experiment with various values [^1] We test $\alpha\in\{0.1,0.2,0.3,0.4,0.5,0.55,0.6,0.65,0.7,0.75,0.8\}$ for language and arithmetic tasks. of $\alpha$ while fixing $\beta=0.1$. For each perturbation level, we perform distillation until the model reaches at least 95% of the teacher model’s performance on the retain set or completes training on one full epoch of pretraining data, whichever comes first. Lower-level experimental details are available in Appendix C.
Figure 5: Unlearning robustness scales with more perturbation. (a, c) UNDO scaling trend for $\alpha$ between 0.1 and 0.8 and $\beta=0.1$, showing trade-off between robustness measured as $(P_{\text{UNDO}}-P_{\text{Unlearn Only}})/(P_{\text{Data Filtering}}-P_{\text{Unlearn Only}})$ where $P$ is performance, and compute measured as $S_{\text{UNDO}}/S_{\text{Data Filtering}}$ where $S$ is training steps. Points denote median values, error bars show variation across five random seeds. (b) Relearning trends for Korean domain with $\alpha=\{0.2,0.4,0.6,0.8\}$. (d) Relearning trends for Multiplication & Division with $\alpha=\{0.55,0.65,0.7,0.75\}$.
Result: UNDO enables trading off compute for robustness to relearning. Figure 5 demonstrates that increasing the perturbation parameter $\alpha$ enhances robustness against relearning attacks while requiring proportionally more computation to recover the teacher model’s performance on the retain set. This relationship appears roughly linear for language unlearning in Figure 5 (a) and exhibits a steeper initial curve for arithmetic unlearning in Figure 5 (c). Figures 5 (b, d) show varying degrees of resistance to relearning across different $\alpha$ values, with higher values consistently demonstrating improved robustness in both domains. For language unlearning, even relatively small perturbations ($\alpha=0.2$) produce noticeable improvements in relearning resistance, while arithmetic unlearning shows meaningful improvements at higher perturbation levels ($\alpha\geq 0.55$).
Figure 6: Effects of varying $\alpha$ and $\beta$ noise parameters. (a-d) Relearning curves for UNDO models with fixed mixing coefficient $\alpha$ = {0.1, 0.3, 0.5, 0.7} while varying noise scale $\beta$ = {0.05, 0.1, 0.5, 1.0}. Higher mixing coefficients ($\alpha$) consistently produce more robust unlearning (flatter curves), while the effect of noise scale ($\beta$) is more nuanced. (e) Computational cost matrix showing distillation steps required to recover within 5% of the teacher’s retain performance across $\alpha$ and $\beta$ combinations.
To learn more about the interplay between mixing coefficient ($\alpha$) and noise scale ($\beta$), we conduct a hyperparameter sweep as reported in Figure 6. Changes in mixing coefficient ($\alpha$) produce a more pronounced effect on robustness than the noise scale ($\beta$), which has a less significant impact. This pattern holds across all settings where models are distilled to within ±$\pm 5\%$ of the teacher’s retain performance. While incorporating standard noise from Xavier initialization can help improve robustness for very low $\alpha$ values ($\alpha=0.1$), we find that for most $\alpha$ values tested, the specific value of $\beta$ had less impact. Interestingly, non-standard $\beta$ values (0.05, 0.1, or 0.5) typically worked similarly well or better than $\beta=1.0$ in our language setup. The computational cost matrix in Figure 6 (e) confirms that $\alpha$ is the primary driver of compute requirements, with higher values often requiring more distillation steps.
6 Comparisons with Other Unlearning Methods
We now compare UNDO against existing robust unlearning methods to evaluate their effectiveness across two key metrics, retain performance and resistance to relearning. The ideal unlearning method would preserve high performance on desired tasks while remaining resistant to adversarial attempts to recover forget capabilities through finetuning. We first compare diverse unlearning methods in our language and arithmetic tasks in Figure 7, subjecting models to increasingly stronger adversarial relearning attacks.
We compare against several approaches that address the robust unlearning problem using different mechanisms to resist relearning attacks. Sharpness-Aware Minimization (SAM) [22] optimizes for flat loss regions where small parameter perturbations maintain similar outputs, making the model less sensitive to fine-tuning. Representation Noising (RepNoise) [59] combines unlearning with noise injection at the representation level, creating interference in the model’s internal representations of forget capabilities. UNDIAL [19] also uses distillation for unlearning but differs from our approach by not applying global parameter corruption techniques like noising. For UNDO, we fix $\alpha$ to 0.6 and $\beta$ to 0.1, and distill from a MaxEnt-suppressed model with varying retain thresholds [^2] We test {0.05, 0.09, 0.13, 0.17, 0.21, 0.25, 0.29, 0.33} retrain thresholds for language and arithmetic tasks. to obtain different points along the retain-forget trade-off curve. Figure 5 suggests that higher $\alpha$ values would yield greater robustness, while lower $\alpha$ values would yield less. We discuss other relevant experimental details in Appendix D.
Figure 7: Comparing unlearning methods across different adversarial strengths. (a, d) Initial performance after unlearning but before adversarial attacks. (b, e) Relearned forget performance after moderate relearning (40 steps). (c, f) Performance after extensive relearning (500 steps).
As Figure 7 shows, while more compute-efficient unlearning methods (GradDiff, MaxEnt) achieve good initial retain-forget trade-offs before relearning attacks, they rapidly degrade under adversarial pressure. In general, we observe that methods designed for robustness (SAM, RepNoise) also show significant performance deterioration when subjected to stronger relearning attacks. In contrast, UNDO with MaxEnt maintains more robust unlearning performance across all explored attack strengths. This creates a new Pareto frontier that approaches the gold standard of data filtering with full retraining but requires less compute and data labeling.
Figure 8: UNDO makes MaxEnt and RMU more resilient to relearning WMDP. Relearning trends averaged across 4 seeds for WMDP-Bio (a, b) and WMDP-Cyber (c, d). The retain/forget performance was measured on MMLU/the WMDP benchmark. Best adversaries are bolded.
Moving beyond our synthetic language and arithmetic unlearning tasks, we also test our methods on the more realistic Weapons of Mass Destruction Proxy (WMDP) benchmark [40]. For this setup, we use Gemma-2-2B [71] rather than custom pretrained models. Following our established methodology, we first apply standard unlearning using either MaxEnt or RMU, then apply UNDO with $\alpha=0.25$ and $\beta=0$. We compare this approach against the other robust unlearning baselines. We evaluate robustness against seven diverse adversarial scenarios involving different data mixtures, question-answer formats, model perturbations, and learning rate variations (See Section D.2).
Figure 8 shows that UNDO consistently improves relearning resilience across all WMDP configurations. These results have immediate practical implications. When model distillation is already planned for efficiency, incorporating unlearning before distillation provides robustness benefits at minimal additional cost. This approach integrates capability removal into existing distillation workflows, providing a practical path to robust unlearning.
Limitations. In our WMDP experiments on a pretrained Gemma-2-2B model, the UNDO model achieves 4.88% lower on MMLU on average. While this places UNDO on the Pareto frontier, other methods achieve a similar trade-off between robust forget and retain performance in this setting, as shown in Figure 11. This contrasts with the prior arithmetic and language setting results, where we see clear gains on the Pareto frontier in Figure 7. We hypothesize this difference comes from the fact that we undertrain in the WMDP setting, distilling on only 0.015% of the size of the original pretraining corpus compared to higher percentages in the other synthetic settings. We discuss this more in Section D.2.7). We expect that model developers would not encounter the same challenges and would have sufficient access to compute, pretraining data, and their proprietary methods to accelerate the process. Distillation is considerably more costly than finetuning-based unlearning methods, but we expect this cost to be justified when robust capability removal is critical.
7 Related Work
Machine unlearning aims to remove the influence of specific training examples from a trained model. This term was first introduced by Cao and Yang [10], which explored adjusting a trained model’s outputs to mimic retraining without the data to be forgotten. Machine unlearning has been a significant research focus for differential privacy, implementing the “right to be forgotten” [60] through frameworks that ensure removal of specific datapoints [37, 3, 27, 44, 74]. The field has evolved into exact unlearning [8, 23, 14, 50] and approximate unlearning [26, 63, 41, 76]. Exact unlearning targets that the unlearned model’s parameters exactly match those of a model retrained from scratch without the data to be forgotten [25], while approximate unlearning relaxes this constraint to require only probabilistically similar output distributions [67].
For LLMs, exact unlearning becomes impractical due to scale and non-convexity [42], leading recent studies to frame unlearning as an optimization problem that approximates the behavior of the retrained model [20, 48, 34, 43, 35, 84, 78]. LLM unlearning typically involves finetuning using objectives like maximizing prediction loss on forget sets [33, 85, 5, 21, 88] or aligning predicted token distributions with target distributions using KL divergence [38, 79, 12, 77]. Recent studies demonstrate that existing LLM unlearning methods achieve behavioral suppression rather than true unlearning [30, 88, 73], remaining vulnerable to finetuning attacks that quickly recover suppressed knowledge [31, 47, 18]. This motivates our student-teacher approach [15] for robust unlearning.
8 Conclusion
There is a disconnect between what a model does and what it could do. Our findings illuminate this gap in two ways: first, by showing that even oracle-matched models retain latent capabilities vulnerable to adversarial elicitation; and second, by demonstrating that distillation discards these capabilities while preserving desired behavior. This insight transforms distillation, a standard practice in LLM development, into a security measure, offering a practical path to robust capability removal.
Acknowledgments
We thank Henrik Marklund for very insightful discussions. We are grateful to Rishub Tamirisa for the guidance in navigating WMDP benchmarking procedures and to Eric Easley for sharing valuable strategies for WMDP dataset cleaning and productive discussions about potential improvements to our method. Vivek Hebbar, Andis Draguns, and Jake Mendel offered helpful comments on the abstract.
We thank Iftekhar Uddin and Laura Vaughan for facilitating access to computational resources and funding support. We thank Lisa Wong for help with the design of Figure 1.
Bruce W. Lee’s work on this project was supported by funding from Open Philanthropy and the Long-Term Future Fund. We also thank MATS for facilitating our collaboration and providing the institutional framework that made this research possible.
Appendix A Experimental Details for Section 3
In Section 3, we present oracle matching experiments. Here, we present the specific experimental details used to obtain the results presented.
Dataset  Our process starts from collecting datasets to facilitate our language and arithmetic unlearning experiments. For language experiments, we utilize two sources: (1) HuggingFaceFW/fineweb-2333https://huggingface.co/datasets/HuggingFaceFW/fineweb-2 Korean subset, providing non-English language examples, and (2) HuggingFaceFW/fineweb-edu [^4] https://huggingface.co/datasets/HuggingFaceFW/fineweb-edu, containing English language examples with high-quality content [57, 56]. We sample 10 million rows from each source.
For arithmetic experiments, we generate 1 million arithmetic examples covering four basic operations: addition, subtraction, multiplication, and division. We control the range of operands, using integers from 1 to 50 for addition and subtraction, and integers from 1 to 20 for multiplication and division. For division problems, we ensure clean division with no remainders by constructing problems where the dividend is a product of the divisor and quotient.
The arithmetic dataset consists of two formats:
  • Equation format: Direct mathematical expressions (e.g., "12 + 7 = 19", "36 / 9 = 4")
  • Word problems: Natural language scenarios generated using templates with randomized variables including names, objects, places, and quantities
Word problems are constructed using templates. For example, addition problems might involve combining collections of objects ("Emma has 8 marbles and Jack has 13 marbles. How many marbles do they have altogether?"). Division problems might involve distributing items ("There are 24 cookies in a container. If each student gets six cookies, how many students will receive cookies?").
Both language and arithmetic datasets have separate test sets. For language data, we allocated 1,000,000 tokens for validation, which consists of 500,000 tokens for the retain domain (English) and 500,000 tokens for the forget domain (Korean). For arithmetic data, we allocated 1,000,000 questions for validation, which consists of 500,000 questions for the retain domain (addition and subtraction) and 500,000 questions for the forget domain (multiplication and division).
Data Processing  All datasets are processed using the Google Gemma-2-2b tokenizer. Documents are chunked to a maximum length of 2,048 tokens, with documents shorter than 50 tokens being filtered out. Processing is done in parallel using multiple CPU cores to tokenize and format the data efficiently. We use a fixed random seed of 42 for data sampling and shuffling to ensure reproducibility.
Model Architecture  We use scaled-down versions of the Gemma-2 architecture for our language and arithmetic unlearning experiments. The model architectures and their key parameters are detailed in Table 1.
Parameter Gemma-2-0.1B Gemma-2-0.3B
Total Parameters $\sim$100M $\sim$300M
Hidden Layers 14 14
Attention Heads 8 8
Hidden Size 320 768
Intermediate Size 1280 3072
Head Dimension 128 128
Table 1: Model Architecture Parameters
The Gemma-2-0.1B model is used for the language unlearning experiments, while the Gemma-2-0.3B model is used for the arithmetic unlearning experiments. We intentionally vary this to show results across different model sizes. We chose these parameters based on the ratio used in the original Gemma-2 architecture to maintain the model’s core properties.
Pretraining  We then pretrain our Oracle and Base models. Table 2 details the hyperparameters used during pretraining. We reference these hyperparameters from [69, 87, 7].
Language Pretraining Arithmetic Pretraining
Hyperparameter Base Oracle Base Oracle
Model Architecture Gemma-2-0.1B Gemma-2-0.1B Gemma-2-0.3B Gemma-2-0.3B
Parameters 100M 100M 300M 300M
Training Data English + Korean English All Ops + English Add/Sub + English
Data Mixture Ratio 1:1 - 1:3 1:7
Learning Rate 4e-4 4e-4 4e-4 4e-4
Min Learning Rate 4e-5 4e-5 4e-4 4e-4
Batch Size 5 5 15 15
Gradient Accum. Steps 80 80 24 24
Weight Decay 0.1 0.1 0 0
Epochs 1 1 1 1
Maximum Steps - - 1000 1000
LR Schedule Cosine Cosine Cosine Cosine
Warmup Steps 50 50 50 50
Max Sequence Length 2048 2048 256 256
Table 2: Pretraining Hyperparameters
For the language models, the Base model is trained on both English and Korean data with equal probabilities, while the Oracle model is trained exclusively on English data. Both models used identical optimization parameters. For the arithmetic models, the Base model is trained on all four arithmetic operations (addition, subtraction, multiplication, division) mixed with English text at a 1:3 ratio. The Oracle model is trained only on addition and subtraction operations mixed with English at a 1:7 ratio. We end up using $\sim 1B$ pretraining tokens for Gemma-2-0.1B (Language, Oracle), $\sim 2B$ tokens for Gemma-2-0.1B (Language, Base), $\sim 100M$ tokens for Gemma-2-0.3B (Arithmetic, Oracle), $\sim 100M$ tokens for Gemma-2-0.3B (Arithmetic, Base).
All models were trained using the AdamW optimizer with cosine learning rate scheduling. We used a shorter sequence length (256 tokens) for arithmetic models compared to language models (2048 tokens) due to the inherently shorter nature of arithmetic problems.
Oracle Matching  We implemented oracle matching using knowledge distillation with forward KL divergence loss [29] on the pretraining dataset, as also explored in Gemma 2 technical report [71]. As illustrated in Figure 2(a), we examined two configurations: (1) distillation from an Oracle teacher to a randomly initialized model and (2) distillation from an Oracle teacher to a pretrained Base model. Table 3 presents the hyperparameters for these distillation procedures.
Hyperparameter Language Distillation Arithmetic Distillation
Teacher Model Gemma-2-0.1B (Oracle) Gemma-2-0.3B (Oracle)
Student Model Gemma-2-0.1B (Base or Rand Init) Gemma-2-0.3B (Base or Rand Init)
Distillation Data English + Korean All Ops + English
Data Mixture Ratio 1:1 1:3
Learning Rate 4e-4 7e-4
Min Learning Rate 4e-5 7e-5
Batch Size 4 15
Gradient Accum. Steps 160 24
Weight Decay 0.1 0
Maximum Steps - 1000
Epochs 1 1
LR Schedule Cosine Cosine
Warmup Steps 50 50
Max Sequence Length 2048 256
Table 3: Oracle Matching Hyperparameters
Relearning  To evaluate the robustness of unlearning methods, we conduct relearning experiments to assess whether unlearned capabilities could be easily recovered. Notably, fine-tuning is often considered a strong form of latent capability elicitation, so it can provide insights into whether knowledge is truly removed from the model or just suppressed during oracle matching. For all relearning experiments, we use the final student model from the oracle matching procedure as the starting point. Training proceeds for 500 steps, with validation performed every five steps to track both retention and recovery of capabilities. Table 4 presents the hyperparameters used for our relearning experiments.
[…]
Hyperparameter Language Relearning Arithmetic Relearning
Relearning Data Korean Mul/Div + English
Data Mixture Ratio - 1:3
Batch Size 8 15
Gradient Accumulation Steps 1 1
Maximum Steps 500 500
Warmup Steps 1 1
Learning Rate [1e-3, 7e-4, 4e-4, 1e-4, 7e-5, 5e-5, 1e-5, 7e-6, 4e-6, 1e-6]
Evaluation Steps 5 5
Max Sequence Length 2048 256
Table 4: Relearning Hyperparameters
Appendix B Experimental Details for Section 4
In Section 4, we report suppress and distill experiments. Many experimental details are similar to what was reported in Appendix A.
Dataset / Data Processing / Model Architecture / Pretraining / Relearning  See Appendix A.
Unlearning Methods  In Figure 3, we report unlearning trends for three methods: GradDiff, MaxEnt, and RMU. To perform a hyperparameter search, we first manually search for the key hyperparameters (see Table 5).
Hyperparameter Language Unlearning Arithmetic Unlearning
GradDiff MaxEnt RMU GradDiff MaxEnt RMU
Forget Data Korean Korean Korean Mult/Div Mult/Div Mult/Div
Retain Data English English English Add/Sub Add/Sub Add/Sub
Batch Size 4 4 4 40 40 40
Max Steps 100 100 500 10 10 500
Max Length 2048 2048 2048 256 256 256
Alpha - - 1200 - - 200
RMU Layers - - 5–11 - - 5–11
End Layer - - 11 - - 11
c - - 6.5 - - 6
Table 5: Suppress Hyperparameters
B.1 Unlearning Learning Rate Sweeps
We do learning rate sweeps across eight learning rates per method. We choose the one that gives the best trade-off between the retain and forget performance. The chosen values (marked with thick lines in Figure 3) are marked in bold below.
GradDiff (Language, CE Loss)
  • 1e-5
    Initial Retain: 4.9359 → Final Retain: 5.0264 || Initial Forget: 3.4668 → Final Forget: 30.5592
  • 2e-5
    Initial Retain: 4.9359 → Final Retain: 4.9924 || Initial Forget: 3.4668 → Final Forget: 31.5783
  • 3e-5
    Initial Retain: 4.9359 → Final Retain: 4.9967 || Initial Forget: 3.4668 → Final Forget: 32.3159
  • 4e-5
    Initial Retain: 4.9359 → Final Retain: 5.0085 || Initial Forget: 3.4668 → Final Forget: 32.9097
  • 5e-5
    Initial Retain: 4.9359 → Final Retain: 5.0216 || Initial Forget: 3.4668 → Final Forget: 33.5308
  • 6e-5
    Initial Retain: 4.9359 → Final Retain: 5.0334 || Initial Forget: 3.4668 → Final Forget: 34.1291
  • 7e-5
    Initial Retain: 4.9359 → Final Retain: 5.0563 || Initial Forget: 3.4668 → Final Forget: 34.685
  • 8e-5
    Initial Retain: 4.9359 → Final Retain: 5.0762 || Initial Forget: 3.4668 → Final Forget: 35.129
GradDiff (Arithmetic, Accuracy)
  • 6e-6
    Initial Retain: 0.875 → Final Retain: 0.9125 || Initial Forget: 0.8675 → Final Forget: 0.8375
  • 7e-6
    Initial Retain: 0.875 → Final Retain: 0.9075 || Initial Forget: 0.8675 → Final Forget: 0.8325
  • 8e-6
    Initial Retain: 0.875 → Final Retain: 0.815 || Initial Forget: 0.8675 → Final Forget: 0.015
  • 9e-6
    Initial Retain: 0.875 → Final Retain: 0.74 || Initial Forget: 0.8675 → Final Forget: 0
  • 1e-5
    Initial Retain: 0.875 → Final Retain: 0.7475 || Initial Forget: 0.8675 → Final Forget: 0
  • 2e-5
    Initial Retain: 0.875 → Final Retain: 0.1825 || Initial Forget: 0.8675 → Final Forget: 0
  • 3e-5
    Initial Retain: 0.875 → Final Retain: 0.8025 || Initial Forget: 0.8675 → Final Forget: 0.28
  • 4e-5
    Initial Retain: 0.875 → Final Retain: 0.79 || Initial Forget: 0.8675 → Final Forget: 0.1775
MaxEnt (Language, CE Loss)
  • 1e-5
    Initial Retain: 4.9359 → Final Retain: 4.9439 || Initial Forget: 3.4672 → Final Forget: 11.3281
  • 2e-5
    Initial Retain: 4.9359 → Final Retain: 4.9397 || Initial Forget: 3.4672 → Final Forget: 11.6977
  • 3e-5
    Initial Retain: 4.9359 → Final Retain: 4.942 || Initial Forget: 3.4672 → Final Forget: 11.9844
  • 4e-5
    Initial Retain: 4.9359 → Final Retain: 4.9466 || Initial Forget: 3.4672 → Final Forget: 12.1715
  • 5e-5
    Initial Retain: 4.9359 → Final Retain: 4.9551 || Initial Forget: 3.4672 → Final Forget: 12.2584
  • 6e-5
    Initial Retain: 4.9359 → Final Retain: 4.9633 || Initial Forget: 3.4672 → Final Forget: 12.3114
  • 7e-5
    Initial Retain: 4.9359 → Final Retain: 4.976 || Initial Forget: 3.4672 → Final Forget: 12.3608
  • 8e-5
    Initial Retain: 4.9359 → Final Retain: 4.9895 || Initial Forget: 3.4672 → Final Forget: 12.3808
MaxEnt (Arithmetic, Accuracy)
  • 6e-5
    Initial Retain: 0.875 → Final Retain: 0.86 || Initial Forget: 0.8675 → Final Forget: 0.1275
  • 7e-5
    Initial Retain: 0.875 → Final Retain: 0.83 || Initial Forget: 0.8675 → Final Forget: 0.0525
  • 8e-5
    Initial Retain: 0.875 → Final Retain: 0.8075 || Initial Forget: 0.8675 → Final Forget: 0.0175
  • 9e-5
    Initial Retain: 0.875 → Final Retain: 0.795 || Initial Forget: 0.8675 → Final Forget: 0.0125
  • 1e-4
    Initial Retain: 0.87 → Final Retain: 0.81 || Initial Forget: 0.87 → Final Forget: 0.035
  • 2e-4
    Initial Retain: 0.87 → Final Retain: 0.4825 || Initial Forget: 0.87 → Final Forget: 0
  • 3e-4
    Initial Retain: 0.87 → Final Retain: 0.045 || Initial Forget: 0.87 → Final Forget: 0.015
  • 4e-4
    Initial Retain: 0.87 → Final Retain: 0.005 || Initial Forget: 0.87 → Final Forget: 0.0225
RMU (Language, CE Loss)
  • 1e-5
    Initial Retain: 4.9359 → Final Retain: 5.4614 || Initial Forget: 3.4668 → Final Forget: 7.2717
  • 2e-5
    Initial Retain: 4.9359 → Final Retain: 6.0976 || Initial Forget: 3.4668 → Final Forget: 8.8457
  • 3e-5
    Initial Retain: 4.9359 → Final Retain: 6.8464 || Initial Forget: 3.4668 → Final Forget: 9.8125
  • 4e-5
    Initial Retain: 4.9359 → Final Retain: 7.6585 || Initial Forget: 3.4668 → Final Forget: 10.2192
  • 5e-5
    Initial Retain: 4.9359 → Final Retain: 8.4008 || Initial Forget: 3.4668 → Final Forget: 10.6792
  • 6e-5
    Initial Retain: 4.9359 → Final Retain: 9.0205 || Initial Forget: 3.4668 → Final Forget: 10.8073
  • 7e-5
    Initial Retain: 4.9359 → Final Retain: 9.4909 || Initial Forget: 3.4668 → Final Forget: 10.958
  • 8e-5
    Initial Retain: 4.9359 → Final Retain: 9.8514 || Initial Forget: 3.4668 → Final Forget: 11.2771
RMU (Arithmetic, Accuracy)
  • 6e-6
    Initial Retain: 0.875 → Final Retain: 0.7575 || Initial Forget: 0.8675 → Final Forget: 0.2275
  • 7e-6
    Initial Retain: 0.875 → Final Retain: 0.68 || Initial Forget: 0.8675 → Final Forget: 0.1375
  • 8e-6
    Initial Retain: 0.875 → Final Retain: 0.615 || Initial Forget: 0.8675 → Final Forget: 0.0675
  • 9e-6
    Initial Retain: 0.875 → Final Retain: 0.52 || Initial Forget: 0.8675 → Final Forget: 0.035
  • 1e-5
    Initial Retain: 0.875 → Final Retain: 0.4425 || Initial Forget: 0.8675 → Final Forget: 0.025
  • 2e-5
    Initial Retain: 0.875 → Final Retain: 0.14 || Initial Forget: 0.8675 → Final Forget: 0
  • 3e-5
    Initial Retain: 0.875 → Final Retain: 0.065 || Initial Forget: 0.8675 → Final Forget: 0
  • 4e-5
    Initial Retain: 0.875 → Final Retain: 0.05 || Initial Forget: 0.8675 → Final Forget: 0
B.2 Distillation
Distillation  In Figure 4, we report relearning trends for the suppressed, distilled, and gold standard. We use the same dataset that was used to pretrain respective models as distillation data. We also ensure that the distilled model undergoes the same number of update steps as its pretrained counterpart. We perform distillation with the hyperparameters in Table 6. Additionally, we report the distillation trend in Figure 9. To find information about Data Filtering (Gold Standard) models, refer to Appendix A.
Parameter Language Distillation Arithmetic Distillation
GradDiff MaxEnt RMU GradDiff MaxEnt RMU
Teacher Model Suppress Only [^5] This refers to the GradDiff/MaxEnt/RMU unlearned models, which are also marked as ”Suppress Only” in Figure 4. Suppress Only Suppress Only Suppress Only Suppress Only Suppress Only
Student Model Random Init Random Init Random Init Random Init Random Init Random Init
Distillation Data Kor + Eng Kor + Eng Kor + Eng All Ops + Eng All Ops + Eng All Ops + Eng
Interleave Probs [0.5, 0.5] [0.5, 0.5] [0.5, 0.5] [0.75, 0.25] [0.75, 0.25] [0.75, 0.25]
Batch Size 4 4 4 15 15 15
Grad. Accum. Steps 60 60 60 24 24 24
Epochs 1 1 1 1 1 1
Learning Rate 9e-4 9e-4 9e-4 7e-4 7e-4 7e-4
Min Learning Rate 7e-4 7e-4 7e-4 7e-5 7e-5 7e-5
Max Steps - - - 1000 1000 1000
Warmup Steps 50 50 50 50 50 50
Scheduler Cosine Cosine Cosine Cosine Cosine Cosine
Weight Decay 0.1 0.1 0.1 0.0 0.0 0.0
Max Seq. Length 2048 2048 2048 256 256 256
Table 6: Distillation Hyperparameters
Figure 9: Distillation Trends
Figure 10: Figure 4 shown with average lines bolded.
Appendix C Experimental Details for Section 5
In Section 5, we introduce UNDO as an approach to leverage random noising as an alternative to full random initialization.
Dataset / Data Processing / Model Architecture / Pretraining / Relearning  See Appendix A.
Unlearning Method  We choose the best MaxEnt method from Appendix B based on its favorable trade-off between retain and forget performance. For language experiments, we use a learning rate of 8e-5, which demonstrated the highest trade-off score while maintaining strong performance on both domains (Initial Retain: 4.9359 → Final Retain: 4.9895, Initial Forget: 3.4672 → Final Forget: 12.3808). For arithmetic experiments, we use a learning rate of 9e-5, which similarly achieved strong performance (Initial Retain: 0.875 → Final Retain: 0.795, Initial Forget: 0.8675 → Final Forget: 0.0125).
Noising Method  The novel contribution introduced in Section 5 is controlled parameter corruption, allowing us to interpolate between the original model and random noise. This provides a more flexible approach compared to binary choices of keeping original parameters or fully reinitializing them. The algorithm applies a weighted combination of the original parameters and scaled random noise based on two hyperparameters: $\alpha$ controls the proportion of corruption applied, while $\beta$ controls the magnitude of the noise. The pseudocode is as follows:
def do_corruption(model, noise_alpha, noise_beta=0.1, seed=42):
    # Loop through all parameters and add random noise
    assert 0 <= noise_alpha <= 1
    assert 0 <= noise_beta
    for param in model.parameters():
        if param.requires_grad:
            # Initialize corruption tensor
            corruption = torch.zeros_like(param.data)
            # Generate appropriate noise based on parameter dimensionality
            if len(param.data.shape) == 2:
                # For weight matrices (2D tensors), use Xavier init
                noise = torch.nn.init.xavier_uniform_(
                    torch.empty_like(param.data)
            elif len(param.data.shape) == 1:
                # For bias vectors (1D tensors), use zeros
                noise = torch.zeros_like(param.data)
            else:
                raise RuntimeError(
                    f"Unsupported parameter shape: {param.data.shape}"
            # Scale the noise by beta
            corruption = noise_beta * noise
            # Apply weighted combination
            param.data=(1 - noise_alpha)*param.data+ noise_alpha*corruption
    # Move model to appropriate device
    model.to(device)
For our experiments, we systematically vary $\alpha$ from 0 to 0.8, while maintaining $\beta$ at a constant value of 0.1 in Figure 5. This allows us to evaluate the effect of gradually increasing levels of parameter corruption on both retention of desired capabilities and forgetting of undesired ones.
Distillation Method  See Appendix B. We follow the same distillation protocol established in our previous experiments, using the noised models (obtained after applying the corruption function) as the student models in the distillation process.
Appendix D Experimental Details for Section 6
In Section 6, we compare UNDO against several other unlearning methods to evaluate their effectiveness in maintaining a balance between retain performance and resistance to relearning. This appendix provides the detailed experimental setup and hyperparameters used for these comparison experiments.
D.1 Language and Arithmetic Unlearning
Dataset / Data Processing / Model Architecture / Pretraining  We use the same datasets, processing methods, model architectures, and pretraining procedures as described in Appendix A. Specifically, we conduct experiments on both the language (English/Korean) task using Gemma-2-0.1B and the arithmetic (addition-subtraction/multiplication-division) task using Gemma-2-0.3B.
Baseline Unlearning Methods  We compare several unlearning approaches:
  • GradDiff: Gradient Difference based unlearning with the forget domain.
  • MaxEnt: Maximum Entropy based unlearning that maximizes the model’s uncertainty on the forget dataset while maintaining performance on the retain dataset.
  • RMU: Representation Manipulation for Unlearning, which modifies internal model representations for selected layers.
  • SAM: Sharpness-Aware Minimization that optimizes for flat loss regions to resist relearning.
  • RepNoise: Representation Noising that combines standard unlearning with noise injection at the representation level.
  • UNDIAL: Unlearning via distillation approach without global model damaging.
Learning Rate Sweeps  For GradDiff, MaxEnt, RMU, and UNDIAL, we conducted extensive learning rate sweeps to find optimal configurations. We tested logarithmically spaced learning rates: 3e-3, 8e-3, 3e-4, 8e-4, 3e-5, 8e-5, 3e-6, 8e-6, 3e-7, 8e-7, 3e-8, and 8e-8. For each method, we chose the learning rate that provided the best trade-off between retain performance and forget degradation.
RepNoise Configurations  For RepNoise experiments, we used the best MaxEnt hyperparameters as the base configuration and then conducted a grid search over the following RepNoise-specific parameters:
Alpha ($\alpha$) Beta ($\beta$)
0.1 0.0001
0.1 0.001
0.5 0.01
0.5 0.1
1.0 0.1
1.0 1.0
2.0 2.0
4.0 4.0
Table 7: RepNoise parameter configurations, where alpha controls the strength of the noise loss and beta controls the ascent loss.
SAM Configurations  For SAM experiments, we also started with the best MaxEnt hyperparameters and performed a parameter sweep over the SAM perturbation radius ($\rho$) parameter:
Rho ($\rho$)
0.001
0.003
0.005
0.01
0.02
0.04
0.07
0.1
Table 8: SAM perturbation radius values explored. The value 0.01 is recommended by the original SAM paper.
UNDO Configuration  For our proposed UNDO method, we fixed the corruption parameters at $\alpha=0.6$ and $\beta=0.1$ based on our findings in Section 5. We used the MaxEnt as the unlearning algorithm before applying random noising and distillation. We also experimented with various retain threshold values {0.05, 0.09, 0.13, 0.17, 0.21, 0.25, 0.29, 0.33} to obtain different points along the retain-forget trade-off curve.
Common Training Parameters  All language experiments used the following common parameters unless otherwise specified for a particular method:
  • Batch size: 4
  • Max sequence length: 2048
  • Scheduler: Cosine
  • Weight decay: 0.0
  • Gradient clipping threshold: 1.0
For arithmetic experiments, we used:
  • Batch size: 40 (GradDiff: 20, RepNoise/SAM variants: 5 with gradient accumulation steps of 8)
  • Max sequence length: 256
  • Scheduler: Cosine
  • Weight decay: 0.0
  • Gradient clipping threshold: 1.0
Method-Specific Parameters  Each unlearning method required specific hyperparameters:
  • GradDiff: Alpha (retain-forget tradeoff) = 1 for language, Alpha (retain-forget tradeoff) = 15 for arithmetic
  • RMU: Layers = [5-11], End layer = 11, Alpha = 1200, c = 6.5 for language; Alpha = 200, c = 6 for arithmetic
Relearning Evaluation  To evaluate robustness against relearning, we fine-tuned each unlearned model on the forget dataset (Korean or multiplication/division) for varying numbers of steps (0, 40, and 500) to simulate increasingly strong adversarial attacks. We monitored both the retain performance (on English or addition/subtraction) and forget performance to assess how well each method maintained the desired knowledge while suppressing the forgotten capabilities.
D.2 WMDP
We apply our method to the WMDP [40] benchmark on both the Cyber and Bio domain and compare to existing unlearning methods.
We use Gemma-2-2b LABEL:team2024gemma as the base model.
D.2.1 Evaluation/Benchmark
For the retain evaluation we use MMLU, evaluating the model with 5-shot on a subset of 40% of the full MMLU evaluation and report accuracy. For the forget evaluation we use WMDP Bio/Cyber depending on the domain use the entire evaluation, using zero-shot prompting.
D.2.2 Datasets
WDMP includes a WMDP forget and retain corpora for each domain. However, it is common practice to use other datasets as well. For example the original WMDP paper introduces RMU [40] using Wikitext [49] as the retain dataset.
Applying unlearning methods using the original forget and retain datasets as well as public auxiliary datasets such as Wikitext proves difficult, especially for the biology domain and methods other than RMU. We hypothesize this is due the datasets possessing numerous differences that could be used to differentiate them. This allows the unlearning methods to achieve low loss on their objectives but fail to generalize to the evaluation. To remedy this, we create a dataset that eliminates most undesired differences and emphasizes the relevant differences in the domains. We extract a question answer dataset from each corpora, Bio retain, Bio forget, Cyber retain, and Cyber forget, by prompting Gemini flash to write questions and answer from text sampled from the corpora. Additionally, we extract questions and answer from English Wikipedia using the same process. The following is an example of a question and answer generated for the Cyber forget dataset.
Question: How does Address Space Layout Randomization (ASLR) complicate shellcode injection?
Answer: Address Space Layout Randomization (ASLR) randomizes the memory addresses of key program areas, making it harder for attackers to predict where shellcode should be placed in memory for successful exploitation.
We experimented with using different datasets and combinations of datasets to optimize the tradeoff between forgetting and maintaining retain performance. Ultimately, we use the forget question-answer dataset for the forget set and a 50/50 mix of the retain question-answer dataset and Wikitext for the retain set.
D.2.3 Baseline Unlearning Methods
We use two baseline unlearning methods, RMU and MaxEnt.
Both RMU and MaxEnt involve optimizing a loss that is the sum of a retain component and a forget component. For WMDP, we parameterize this combination with $\alpha$ such that $loss=\alpha*retain\_component+(1-\alpha)*forget_{c}omponennt$ where $10<\alpha<$ 1.
For MaxEnt, we use a variation in which the loss maximizes entropy on the forget set while minimizing kl divergence between the base model and unlearned model on the retain set. For RMU, we use the same loss as language/arithmetic settings. We use the parameters in Table 9.
Parameter Value
General Training Parameters
Batch size 40
Sequence length 256
Steps 90
Learning rate Fixed (no scheduler)
Weight decay 0
Warmup steps 0
RMU-Specific Parameters
$c$ value 80
RMU layers $\{10,11,12,13,14,15\}$
RMU end layer 15
Table 9: Hyperparameter Settings
We run hyperparameters sweeps on alpha and learning rate for both methods, then chose three points across the frontier to use in the experiment. We run four seeds of each method and apply adversaries to each to measure robust forget performance. Of the three points we use the one with the highest retain performance as the teacher for the UNDO models. Ultimately, the models in Table 10 are chosen.
Domain Method Configuration Retain Performance Ranking
Bio RMU $\alpha=0.5$, $\text{lr}=5\times 10^{-5}$ Best
$\alpha=0.3$, $\text{lr}=1\times 10^{-4}$ Medium
$\alpha=0.1$, $\text{lr}=1\times 10^{-4}$ Worst
MaxEnt $\alpha=0.3$, $\text{lr}=2\times 10^{-5}$ Best
$\alpha=0.1$, $\text{lr}=2\times 10^{-5}$ Medium
$\alpha=0.75$, $\text{lr}=5\times 10^{-5}$ Worst
Cyber RMU $\alpha=0.5$, $\text{lr}=2\times 10^{-5}$ Best
$\alpha=0.2$, $\text{lr}=2\times 10^{-5}$ Medium
$\alpha=0.1$, $\text{lr}=2\times 10^{-5}$ Worst
MaxEnt $\alpha=0.2$, $\text{lr}=2\times 10^{-5}$ Best
$\alpha=0.3$, $\text{lr}=3.5\times 10^{-5}$ Medium
$\alpha=0.3$, $\text{lr}=5\times 10^{-5}$ Worst
Table 10: Selected Models Ordered by Performance Trade-off
D.2.4 Distillation/UNDO
We apply UNDO using an auxiliary dataset. The dataset consists of the following datasets, specified by the Huggingface name, and mixed at the percentages shown in Table 11.
Dataset Proportion
HuggingFaceFW/fineweb-edu (subset_name="sample-10BT") 0.35
legacy-datasets/wikipedia (subset_name="20220301.en") 0.35
Magpie-Align/Magpie-Llama-3.3-Pro-1M-v0.1 0.05
Magpie-Align/Magpie-Llama-3.1-Pro-1M-v0.1 0.04
Magpie-Align/Llama-3-Magpie-Pro-1M-v0.1 0.04
Magpie-Align/Magpie-Gemma2-Pro-534K-v0.1 0.04
Magpie-Align/Magpie-Phi3-Pro-1M-v0.1 0.04
Magpie-Align/Magpie-Qwen2.5-Pro-1M-v0.1 0.04
Magpie-Align/Magpie-Qwen2-Pro-1M-v0.1 0.04
Wikipedia question-answer 0.01
Total 1.00
Table 11: Auxiliary Dataset Composition
D.2.5 Relearning Adversaries
We extensively explore relearning adversaries via small batches of relearning runs. To ensure fairness and avoid only selecting adversaries that do well on baseline methods, we always run configurations on the UNDO method. We run the most promising configurations across all seeds of the methods, and we run all configurations on UNDO. We report the metrics from the best adversary for each method.
We explore seven configurations and vary the learning rate for each.
We have a variety of techniques and datasets that we vary to form 6 adversary configurations.
Dataset Descriptions:
  1. 1.
    forget/retain: Consists of a 50/50 mixed forget and retain WMDP corpora from the corresponding domain.
  2. 2.
    forget/retain-qa: Consists of a 50/50 mixed forget and retain question-answer dataset from the corresponding domain.
  3. 3.
    wiki-qa: Consists of a 50/25/25 mix of wikitext, forget question-answer, and retain question-answer from the corresponding domain.
Sequence Length Descriptions:
  • Standard: Sequence length 256
  • Long: Sequence length 1024
Technique:
  • Shrink-perturb: When applied, we apply shrink perturb to the weights as formulated in the UNDO method with $\alpha=0.05$ before relearning.
Adversary Name Dataset Sequence Length Shrink-perturb
forget/retain 1 standard no
forget/retain-qa 2 standard no
forget/retain-long 1 long no
forget/retain-qa-long 2 long no
wiki-qa-long 3 long no
sp-forget/retain 1 standard yes
sp-forget/retain-qa 2 standard yes
sp-bio-wiki-qa-long 3 long yes
Table 12: Configuration of the seven adversary techniques used to evaluate unlearning robustness. Dataset values correspond to: (1) original forget/retain data, (2) question-answer formatted data, and (3) Wikipedia data with QA formatting. Sequence length indicates whether a standard or extended context was used. Shrink-perturb indicates whether the model parameters were perturbed before relearning.
The diversity of adversaries highlights how each existing method may be robust against some adversaries, but does not deeply remove knowledge and therefore is not robust against all adversaries. For example, repnoise was much more susceptible to the shrink-perturb relearning, while being more robust to other adversaries. Similarly, for the TAR method, the authors report robustness across most adversaries, but select adversaries were completely able to recover performance LABEL:tamirisa2024tamper. While we were unable to find hyperparameters that maintained reasonable retention performance in the time we spent, we expect this to hold true for TAR in all of our settings as well.
D.2.6 Baseline Robust Methods
We explore two robust baselines, SAM and RepNoise, in the Arithmetic and Language settings. For SAM, we set $\rho$ to 0.01, and for RepNoise we use $\beta$=1.0 and $\alpha$=1.0. Both of these methods operate in addition to MaxEnt, so we use the top two selected MaxEnt hyperparameters.
D.2.7 WMDP Pareto Frontier and Discussion
Figure 11: We measure and plot forget and retain accuracy on WMDP Bio (col 1) and Cyber (col 2), immediately after unlearning (row 1), after a low compute adversarial retraining (row two), and after a high compute adversarial retraining (row 3).
We apply MaxEnt, RMU, UNDO, RepNoise, and SAM to Gemma-2-2b. We observe UNDO is on the Pareto frontier for WMDP-Bio, and competitive with other methods for WMDP-Cyber, despite decreases in retain performance. The results are shown in Figure 11. We observe that the UNDO method has a similar robust-forget vs retain tradeoff to other baseline methods.
This contrasts with the earlier results for the language/arithmetic settings seen in Figure 7, with UNDO far exceeding the Pareto frontier set by the baseline methods. We hypothesize that this is due to under-training the model due to limited compute and not having access to the original pretraining corpus.
More thoroughly, the key difference is in the arithmetic/language pareto Figure 7 we use $\alpha=0.6$, whereas in WMDP pareto Figure 11, we use $\alpha=0.25$, implying more noise in the language/arithmetic settings than WMDP. Higher alpha damage to the model’s capabilities, making them harder to retrain. We show this in Figure 5, where higher alpha values are more robust and achieve lower retrained forget performance. However, when we apply higher alpha values to Gemma-2-2b, we find that it doesn’t recover retain performance (see Table 13).
In Figure 5, we observe that higher alpha values require more data to recover performance. For $\alpha=0.6$, the language and arithmetic settings require around 35% and 20% of their pretraining corpus, respectively. For WMDP, our reference model, Gemma-2-2b was originally trained on 2 trillion tokens [70], but we distill on only 300 million, 0.015% of the Gemma-2-2b pretraining corpus. However, for AI companies aiming to apply our method to their models, we expect that they have sufficient compute resources to train on a larger fraction of the pretraining corpus. Additionally, it’s possible that the original pretraining corpus has better coverage or diversity to effectively learn the capabilities compared to our distillation data.
We expect that the challenges we face would not apply to the companies that could apply our methods to their models because they would have sufficient compute resources, access to the full pretraining corpora, and could even have proprietary distillation methods that allow faster distillation or better generalization compared to our setup.
Alpha Retrained MMLU Value
Original (0) 0.55
0.25 0.49
0.35 0.37
0.6 0.25
Table 13: Alpha noise values applied to Gemma-2-2b and corresponding retrained MMLU values. $\beta$ = 0 for all experiments.
Appendix E Compute requirements
We run all experiments on servers with multiple H200 or A100 GPUs. Language and arithmetic pretraining took 4xH200/1xA100 GPUs for two to three days. All distillation and UNDO runs for language are run on 4xH200 GPUs, while relearning took one H200 GPU per setup. All distillation and UNDO runs for arithmetic are run on a single A100 GPU, while relearning took one A100 GPU per setup. Our most computationally expensive experiments are WMDP UNDO runs, which take around 7 hours on one H200. Additionally, computing evaluations while relearning takes 8 minutes to do a sparse evaluation relearning run (which collects evaluations at steps 10, 25, 50, 150, and 500).
Appendix F Author Contribution
All authors contributed to the early development of the ideas of the paper.
Bruce W. Lee was a core contributor to our experimental codebase and was responsible for framing and conducting most experiments reported in the paper, except WMDP. This includes Unlearn-and-Distill and UNDO experiments reported in Figures 1, 2, 3, 4, 5, 6, 7. BL implemented baselines and plotted all figures in the paper. BL wrote the majority of the later sections and assisted in revising the earlier sections.
Addie Foote conducted all WMDP experiments in Figures 8, 11 and initially designed our arithmetic unlearning experiments. AF was a core contributor to our experimental codebase, making unlearning and distillation scripts, efficiency optimizations, and experiment infrastructure. AF wrote key sections on the WMDP experimental setup and the limitations discussion. AF consistently addressed the most challenging and uncertain problems, operating in the face of uncertainty, particularly with the WMDP benchmark.
Alex Infanger contributed significantly to the writing of the paper, the development of the problem statement, the interpretation and the discussion of experimental results, and the implementation of a pilot example of the oracle matching result. AI embarked on a semantic odyssey to improve our understanding of the robust unlearning problem.
Leni Shor conducted exploratory research to mechanistically validate our oracle matching experiments, helping the team gain confidence in our ideas. LS also explored formalisms that improved our understanding of the problem we were trying to solve. LS conducted initial MNIST experiments with AI, streamlining the MNIST codebase.
Harish Kamath achieved the first working implementation of our experiments on WMDP, which informed our later experimental approach. HK implemented and conducted experiments with noising schedulers as part of the initial UNDO framework development, and explored activation-based distillation instead of logit-based distillation.
Jacob Goldman-Wetzler developed the initial proof-of-concept demonstrating that distillation robustifies unlearning and reported the first positive results. JGW assisted with WMDP experiments and optimized RMU on Gemma models. JGW also contributed to activation-based distillation experiments and provided debugging support for the codebase when needed.
Bryce Woodworth served as research manager, organizing and facilitating team research meetings and coordinating project logistics. BW provided essential support in maintaining team communication and collaboration throughout the research process. BW contributed to copy-editing and offered valuable feedback on presentation and clarity.
Alex Cloud suggested combining distillation with unlearning and investigating oracle matching, and also co-led the research team. AC worked closely with the team, offering helpful input on implementation details, experimental design, and conceptual framing. He also contributed to writing.
Alexander Matt Turner co-led the research team and was instrumental in the initial conceptualization of the project. AT contributed to earlier iterations of the distillation approach, originally explored in the context of Gradient Routing based on his suggestions. Throughout the project, AT provided valuable feedback on experimental direction, helped analyze results, and contributed to manuscript development.